OpenStudy (anonymous):

if cos (pie x)=x^2-x-5/4 then the value of x is

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

cos pie x=(4x^2-4x-5)/4

OpenStudy (anonymous):

2<=(2x^2-1)<=10

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Firstly, I am really \(\color{green}{\textsf{Sorry}}\).. I was busy at some another post.. :(

OpenStudy (anonymous):

\[\cos (\pi x)=x^2-x-\frac{5}{4}\]

OpenStudy (anonymous):

Is this the question??

OpenStudy (anonymous):

@rosy2 are you there?? Are you alive?? :P

OpenStudy (gorv):

cos will give values between -1 and 1

OpenStudy (gorv):

-1<=(4x^2-4x-5)/4<=1 multiply both side by 4

OpenStudy (gorv):

-4<(4x^2-4x-5)<4

OpenStudy (gorv):

let forget equal to sign we will add it at the end

OpenStudy (gorv):

now we have 2 equation -4<4x^2-4x-5 and 4x^2-4x-5<4

OpenStudy (anonymous):

yes @waterineyes ,@gorv i was working on some other problem.

OpenStudy (gorv):

u got upto here??

OpenStudy (anonymous):

yes

OpenStudy (gorv):

@rosy2

OpenStudy (gorv):

ok can u solve the inequality ??

OpenStudy (gorv):

-4<4x^2-4x-5 add 4 on both side

OpenStudy (gorv):

4x^2-4x-5+4>0

OpenStudy (gorv):

4x^2-4x-1>0

OpenStudy (gorv):

4x^2-4x-5<4 subtract 4 from both side

OpenStudy (gorv):

4x^2-4x-9>0

OpenStudy (gorv):

we have to solve these 2 equation

OpenStudy (anonymous):

ok i will work and say the solution

OpenStudy (gorv):

okk..u try n if get a prob we will help u thn

OpenStudy (anonymous):

1/2,(1+or-16sqrt(2))/2

OpenStudy (gorv):

we have to find where they are greater than zero or less than zero

OpenStudy (gorv):

i dont have calc with me right now

OpenStudy (anonymous):

is it correct?

OpenStudy (gorv):

can i tell u the ans in the eve ??

OpenStudy (anonymous):

okay

OpenStudy (gorv):

4x^2-4x-1>0 \[\frac{ 4\pm \sqrt{(-4)^2-4*4*(-1)} }{ 2*4 }\]

OpenStudy (gorv):

\[\frac{ 4 \pm \sqrt{16+16}}{ 8}\]

OpenStudy (gorv):

\[\frac{ 4 \pm \sqrt{16*2}}{ 8 }=\frac{ 4\pm4\sqrt{2} }{ 8 }\]

OpenStudy (gorv):

\[\frac{ 1 \pm \sqrt{2}}{ 2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(4\pm \sqrt{160})/8\]

OpenStudy (anonymous):

\[(1\pm4\sqrt{10})/2\]

OpenStudy (gorv):

very veryy good rosy

OpenStudy (anonymous):

\[(1\pm4\sqrt{10})/2\]

OpenStudy (gorv):

u did a little mistake

OpenStudy (gorv):

4 will not come before sqrt

OpenStudy (gorv):

\[\frac{ 4 }{ 8} \pm \frac{ 4\sqrt{10} }{ 8 }=\frac{ 1 \pm \sqrt{10}}{ 2 }\]

OpenStudy (gorv):

got me??

OpenStudy (anonymous):

-9*16=144 add it to 16 we get160 right?